OpenStudy (anonymous):

Solve the following equation for s 6s=p

OpenStudy (anonymous):

6s=p in order to get s alone on the one side, you have to DIVIDE both sides by six (because the s is multiplied by 6 Therefore, 6s/6=p/6 Simplify from there and you'll get s=p/6

OpenStudy (anonymous):

so s = p/6?

OpenStudy (anonymous):

yes